OpenStudy (erniellyd28):

add the polynomials -9X To the second +4x 2x To the second _ 8x-6

TheSmartOne (thesmartone):

Btw, it's not: to the second, it's squared so we have -9x^2 + 4x being added to 2x^2 - 8x - 6 let's put them side by side -9x^2 + 4x + 2x^2 - 8x - 6 can you add the like terms? -9x^2 + 2x^2 = ? 4x - 8x = ?

OpenStudy (erniellyd28):

84+-4?

TheSmartOne (thesmartone):

no, remember: -9x^2 + 2x^2 = (-9 + 2)x^2 = ?? 4x - 8x = (4 - 8)x = ??